what number and its square differ by 20?
Answer by algebrahouse.com(1659)

"what number and its square differ by 20?"

x = the number

x^2 - x = 20 {what number(x) and its square(x^2) differ by 20}
x^2 - x - 20 = 0 {subtracted 20 from both sides}
(x - 5)(x + 4) = 0 {factored into two binomials}
x - 5 = 0 or x + 4 = 0 {set each factor equal to 0}
x = 5 or x = -4 {solved each equation for x}

The number could be 5 or -4
www.algebrahouse.com